OpenStudy (anonymous):

Solve the following system by elimination. 3x - 2y = 18 5x + 2y = 14 A. (0, 7) B. (4, -3) C. (6, -6) D. (2, 2)

OpenStudy (anonymous):

B. (4, -3)

OpenStudy (anonymous):

3x-2y=18_5x+2y=14 Add the two equations together to eliminate y from the system. 5x+2y=14_<U>3x-2y=18<u>_8x =32 Divide each term in the equation by 8. x=4 Substitute the value found for x into the original equation to solve for y. 3(4)-2y=18 Multiply 3 by each term inside the parentheses. 12-2y=18 Move all terms not containing y to the right-hand side of the equation. -2y=6 Divide each term in the equation by -2. y=-3 This is the final solution to the independent system of equations. x=4_y=-3

OpenStudy (anonymous):

thanks